0.010 M NH4OH (1.0%)
media%2Fba3%2Fba3cc5f6-d7dd-4ba3-a2c4-5b
Answer #1


Degree of dissociation(x) = 1/100 = 0.01

C = concentration of NH4OH = 0.01 M

NH4OH - WEAK base

[OH-] = CX

      = 0.01*0.01

      = 1*10^-4 M

pOH = -log(OH-)

     = -log(1*10^-4)

     = 4

pH = 14-pOH

    = 14-4 = 10

[H3O+] = 10^-10